OpenStudy (anonymous):

21 root beer, 2 cream soda, and 2 apple. you reach into the bag to pull out 1 piece. find the probability that the piece he pulls out is not apple flavored?

OpenStudy (anonymous):

1 - P(apple). This is the complement of getting an apple piece. What's the probability of getting an apple piece?

OpenStudy (anonymous):

is it 1/25? because there are 25 in all so if he pulls out one it'll be 24/1

OpenStudy (anonymous):

P(Apple) = 2/25, thus 1 - P(Apple) = 1 - 2/25 = 23/25